# Work extraction with feedback control using limited resources

### v1

@@ III.3 The “cheeseburger effect”
To understand this phenomenon, we first note that a cheeseburger’s macrostate corresponds to a distribution of microstates representing the distinct arrangements of molecular ingredients composing that cheeseburger. We require two such distributions, corresponding to the two possible macrostates; we denote these distributions $\rho_{L}(\mathbf{x})$ for occupancy of the left box, and $\rho_{R}(\mathbf{x})$ for occupancy of the right box, with argument $\mathbf{x}$ representing a specific molecular arrangement. In the absence of information about in which box the cheeseburger is placed, its distribution would be ${\rho}_{0}=[\rho_{L}(\mathbf{x})+\rho_{R}(\mathbf{x})]/2$.
@@
We now suppose that the macroscopic position of the cheesburger is accessible to measurement (e.g., by opening the boxes to see which one contains the cheeseburger), so that we identify the space $\mathbb{M}$ as consisting only of two possible discrete outcomes, left $m=L$ or right $m=R$, occurring with probabilities $p(L)=p(R)=1/2$. Then by observing, say, $m=L$, the updated state of the system would be $\rho_{X|L}(\mathbf{x})=\rho_{L}(\mathbf{x})$, and we may enact the ideal quenching Hamiltonian, $H_{L}(\mathbf{x})=-k_{B}T\log\rho_{L}(\mathbf{x})$. Similarly, if we obtain $m=R$, the state of the system gets updated to $\rho_{X|R}(\mathbf{x})=\rho_{R}(\mathbf{x})$ and the quench $H_{R}(\mathbf{x})=-k_{B}T\log\rho_{R}(\mathbf{x})$ is performed. Notice that, unlike a measurement of the microstate (a distribution over $\mathbb{X}^{d_{S}}$), this is a measurement of the macrostate (a distribution over $\{L,R\}$). This is similar to the case of Szilard’s engine, where the measurement reveals which side of the box the particle is contained in (but not the specific location and/or momentum of the particle).
@@
Let’s then assume that the cheeseburger is in one of the two possible (Bayesian posterior) states $\rho_{X|{m}}$, for $m=\{L,R\}$. If the outcome $m$ is accessible to the feedback controller a work $W_{m}(m)$ as defined in Eq. (8) would be extracted, where we used that in this case the optimal choice of quench in Eq. (14) is just $n^{\ast}(m)=m$. We assume that accessing the cheeseburger will allow for a positive work extraction, $W_{m}(m)>0$ for both $m=L,R$. The average work that the controller would be able to extract is then $\langle W_{m}\rangle=\sum_{m}p(m)W_{m}(m)>0$. In contrast, if the controller has to randomly choose some action $n$ randomly, i.e. without access to the outcome $m$, the amount of extractable work will be reduced. In particular, let’s first consider the situation in which even if the quenching Hamiltonian is chosen at random, $q=\{L,R\}$, with arbitrary probabilities $r(q)$, the true measurement outcome $m$ is going to be revealed to the experimenter just before action, so that the experimenter has the opportunity to halt the quench unless it is correct. Then when we average over measurement results, the work extracted is
$\langle W\rangle_{r}$ $=\sum_{m}p(m)\sum_{q}r(q)W_{m}(m)\delta_{m,q}$ (23)
$=\sum_{m}p(m)r(m)W_{m}(m)>0.$
@@
$\langle\Delta W\rangle_{r}=\sum_{{m}}p({m})W_{m}({m})[1-r(m)]>0.$ (24)
This difference is strictly positive since $r(m)\leq 1$ and $\sum_{m}r(m)=1$. It can be even macroscopically large if the initial system is far enough from thermal equilibrium, so that $W_{m}(m)$ is large on macroscopic scales for all $m$. That is, if the cheeseburger contains enough processable energy, typically $W_{m}\simeq 300{\rm cal}$ for a standard cheeseburger.
An interesting variant of this analysis arises if we cannot halt an incorrect quench. In that case energetic penalties are incurred as a consequence of implementing a wrong quench, which may involve both the work cost of implementing the quench itself [see also Eq. (20)] (the energy needed to pick up and eat the cheeseburger), and also the free energy lost in the subsequent relaxation to equilibrium (the putrefaction of the cheeseburger in the inaccessible box).
@@
$\langle\tilde{W}\rangle_{r}=\langle W\rangle_{r}-\sum_{m}p({m})\sum_{q\neq m}r(q)W_{q}(m),$ (25)
where the second term in the first line represents the total work lost due to the implementation of the quench $H_{k}$ when the state of the system was $\rho_{X|{m}}$ (instead of $\rho_{X|k}$) and $W_{q}(m)=W_{m}(m)-k_{B}TD(\rho_{X|m}||\rho_{X|q})$ which in general can be either positive or negative.
@@
$\langle\Delta\bar{W}\rangle_{r}=$ $\langle\Delta W\rangle_{r}-k_{B}TI(\mathbf{x}:m)$ (26)
@@
which is consistent with Eq. (III.2) for noise-free observation of the macroscopic variable $m=n$ and no limitations in the available quenches, i.e., $H_{n}=-\ln\rho_{X|n}$. Note that in the above expression we have incorporated the erasure cost, $k_{B}TI(\mathbf{x}:m)=k_{B}T(S(\rho_{0})-S(\rho_{X|L})/2-S(\rho_{X|R})/2)=k_{B}TS(p(m))=k_{B}T\log 2$. This is because, in contrast to the feedback control case, random quenching without observation requires no processing of information and hence a fair comparison requires taking that cost into account. The above equation tells us that despite having learned only one bit about $m$, the thermodynamic consequences for the distribution of the microstates $\mathbf{x}\in\mathbb{X}$ can be arbitrarily large.
@@ IV Optimal work extraction with control limitations
Importantly, rather than taking an arbitrary collection of $N$ Hamiltonians, in many situations we would be interested in constructing an optimal set of quenching Hamiltonians $\mathcal{H}$ with a limited dimension $N=|\mathcal{H}|$, given an initial distribution $\rho_{0}(\mathbf{x})$ and given the conditional probability $P(\mathbf{m}|\mathbf{x})$ characterizing the noisy measurement of the system.
As mentioned before, the measurement results $\mathbf{m}$ can be generically distributed into $N$ disjoint but otherwise arbitrary sets $l(n)$ with $n=1,...,N$ to which one of the possible quenching Hamiltonians $H_{n}(\mathbf{x})\in\mathcal{H}$ is attached by using the lookup table defined by the minimization in Eq. (14). Constructing the optimal set $\mathcal{H}$ requires that the $H_{n}$ within each set $l(n)$ minimizes the expected entropy production across all such partition (see Fig. 1). The optimal form of $H_{n}(\mathbf{x})$ given a partition =$\mathcal{L}=\{\ell(n)\}_{n=1}^{N}$ is derived in Appendix A and reads:
@@
Figure 1: Illustration of the construction of an optimal quenching repertoire $\mathcal{H}$. We group all possible post-measurement distributions $\rho_{X|\mathbf{m}}$ in the simplex $\Delta(\mathds{X}^{d_{S}})$ (blue dots) in $N$ sets according to a given partition of the possible measurement results $\mathbf{m}$ in the space $\mathds{M}^{d_{M}}$. The optimal quench $\rho_{n}^{\rm opt}$ (black dot) withing a given set $l(n)$ (green shaded area) is chosen by finding the distribution that minimizes the average entropy production (Kullback-Leibler divergence) when relaxing from all other post-measurement states within that set (purple dots).
Given the optimal form of the elements of the set $\mathcal{H}$ in Eq. (27), we can now compute the average entropy production (or work dissipated) in Eq. (12) for a given choice of partition $\mathcal{L}$:
@@
$=\sum_{n}p(n)\left(S(\rho_{n}^{\rm opt})-\sum_{\mathbf{m}\in l(n)}p(\mathbf{m}|n)S(\rho_{X|\mathbf{m}})\right),$
@@
We note that in Eq. (29), the optimal quenching distribution $\rho_{n}^{\mathrm{opt}}$ is a mixture distribution of $\rho_{X|\mathbf{m}}$ across $\mathbf{m}$ weighted by $p(\mathbf{m}|n)$ (c.f. Eq. (27)), and that the latter parenthesized term in Eq. (29) is the mean entropy of $\rho_{X|\mathbf{m}}$ over the same mixing distribution $p(\mathbf{m}|n)$. As a consequence it may be rewritten as
@@
with $J_{\mathbf{p}}(\mathbf{P}):=S(Q)-\sum_{i}p_{i}S(P_{i})$ denoting the generalized Jensen-Shannon divergence among a weighted collection of distributions $\mathbf{P}=\{P_{i}\}_{i}$ with (normalized) weights $\mathbf{p}=\{p_{i}\}_{i}$. Here $Q(x)=\sum_{i}p_{i}P_{i}(x)$ denotes the mixture distribution of $\{P_{i}(x)\}_{i}$ with weights $\{p_{i}\}_{i}$. (Note, $\sum_{i}p_{i}=1$ and $\sum_{x}P_{i}(x)=1$.) The above result in Eq. (30) implies that the optimal partition is one that minimizes the generalized JSD among the distributions grouped into a common partition block $\ell(n)$, averaged over the action $n$ taken.
@@ V.1 Feedback control with memory limitations and noise-free measurements
As a standard scenario of general interest in stochastic thermodynamics we consider in the following the case in which the measurement results $m_{i}\in\mathbf{m}$ are associated one to one to the system variables $x_{i}\in\mathbf{x}$ and hence $\mathbb{M}=\mathbb{X}$ and $d_{S}=d_{M}$. Moreover, even if we are allowed to optimally choose the set of quenching Hamiltonians in the set $\mathcal{H}$, we assume operation with a finite memory where the number of quenches $N$ is generically much smaller than the state space of the system $N<|\mathbb{X}^{d_{S}}|$, and hence likewise smaller than the possible measurement results $|\mathbb{M}^{d_{M}}|=|\mathbb{X}^{d_{S}}|$.
@@
where $\mathds{1}\{\mathbf{x}\in\ell(n)\}$ is the indicator function within the partition $l(n)$, and $p^{\rm cg}_{0}(n):=\sum_{x\in\ell(n)}\rho_{0}(\mathbf{x})$ is the coarse-grained version of $\rho_{0}(\mathbf{x})$ according to partition $\ell(n)$. That is, the above quenching Hamiltonian takes finite values in the interval $\mathbf{x}\in l(n)$ and it is infinite otherwise.
@@
$\langle S_{\mathrm{tot}}\rangle$ $=\sum_{\mathbf{m}}\rho_{0}(\mathbf{m})D(\delta_{\mathbf{m},\mathbf{x}}||e^{-\beta H_{\nu}})$

### v2

@@ III.3 The “cheeseburger effect”
To understand this phenomenon, we first note that a cheeseburger’s macrostate corresponds to a distribution of microstates representing the distinct arrangements of molecular ingredients composing that cheeseburger. We require two such distributions, corresponding to the two possible macrostates; we denote these distributions $\rho_{A}(\mathbf{x})$ for occupancy of the left box, and $\rho_{B}(\mathbf{x})$ for occupancy of the right box, with argument $\mathbf{x}$ representing a specific molecular arrangement. In the absence of information about in which box the cheeseburger is placed, its distribution would be ${\rho}_{0}(\mathbb{x})=[\rho_{A}(\mathbf{x})+\rho_{B}(\mathbf{x})]/2$.
@@
We now suppose that the macroscopic position of the cheesburger is accessible to measurement (e.g., by opening the boxes to see which one contains the cheeseburger), so that we identify the space $\mathbb{M}$ as consisting only of two possible discrete outcomes, left $m=A$ or right $m=B$, occurring with probabilities $p(A)=p(B)=1/2$. Then by observing, say, $m=A$, the updated state of the system would be $\rho_{X|A}(\mathbf{x})=\rho_{A}(\mathbf{x})$, and we may enact the ideal quenching Hamiltonian, $H_{A}(\mathbf{x})=-k_{B}T\log\rho_{A}(\mathbf{x})$. Similarly, if we obtain $m=B$, the state of the system gets updated to $\rho_{X|B}(\mathbf{x})=\rho_{B}(\mathbf{x})$ and the quench $H_{B}(\mathbf{x})=-k_{B}T\log\rho_{B}(\mathbf{x})$ is performed. Notice that, unlike a measurement of the microstate (a distribution over $\mathbb{X}^{d_{X}}$), this is a measurement of the macrostate (a distribution over $\{A,B\}$). This is a similar measurement than in the case of Szilard’s engine, where the measurement reveals which side of the box the particle is contained in but not the specific location and momentum of the particle.
@@
Let’s then assume that the cheeseburger is in one of the two possible (Bayesian posterior) states $\rho_{X|{m}}$, for $m=\{A,B\}$. If the outcome $m$ is accessible to the feedback controller a work $W(m)$ as defined in Eq. (8) would be extracted, where in this case the optimal choice of quench in Eq. (14) is just $n^{\ast}(m)=m$. We assume that accessing the cheeseburger will allow for a positive work extraction, $W(m)>0$ for both $m=A,B$. The average work that the controller would be able to extract is then $\langle W\rangle=\sum_{m}p(m)W(m)>0$. In contrast, if the controller has to randomly choose some action randomly, i.e. without access to the outcome $m$, the amount of extractable work will be reduced. In particular, let’s first consider the situation in which even if the quenching Hamiltonian is chosen at random, $q=\{A,B\}$, with arbitrary probabilities $r(q)$, the true measurement outcome $m$ is going to be revealed to the experimenter just before action, so that the experimenter has the opportunity to halt the quench unless it is correct. Then when we average over measurement results, the work extracted is
$\langle W^{\prime}\rangle_{r}$ $=\sum_{m}p(m)\sum_{q}r(q)W(m)~{}\delta_{m,q}$ (21)
$=\sum_{m}p(m)r(m)W(m)>0.$
@@
$\langle\Delta W^{\prime}\rangle_{r}=\sum_{{m}}p({m})W({m})\left[1-r(m)\right]>0.$ (22)
This difference is strictly positive since $r(m)\leq 1$ and $\sum_{m}r(m)=1$. It can be even macroscopically large if the initial system is far enough from thermal equilibrium, so that $W(m)$ is large on macroscopic scales for all $m$. That is, if the cheeseburger contains enough processable energy, typically $W(m)\simeq 300{\rm cal}$ for a standard cheeseburger.
An interesting variant of this analysis arises if we cannot halt an incorrect quench. In that case energetic penalties are incurred as a consequence of implementing a wrong quench, which may involve both the work cost of implementing the quench itself [see Eq. (18)] (the energy needed to pick up and eat the cheeseburger), and also the free energy lost in the subsequent relaxation to equilibrium (the putrefaction of the cheeseburger in the inaccessible box).
@@
$\langle{W}\rangle_{r}=\langle W^{\prime}\rangle_{r}-\sum_{m}p({m})\sum_{q\neq m}r(q)W_{q}(m),$ (23)
where the second term in the first line represents the total work lost due to the implementation of the quench $H_{k}$ when the state of the system was $\rho_{X|{m}}$ (instead of $\rho_{X|k}$) and $W_{q}(m)=W(m)-k_{B}TD(\rho_{X|m}||\rho_{X|q})$ which in general can be either positive or negative.
@@
$\langle\Delta{W}\rangle_{r}=$ $\langle\Delta W^{\prime}\rangle_{r}-k_{B}TI_{X;M}$ (24)
@@
which is consistent with Eq. (III.2) for noise-free observation of the macroscopic variable $m=n$ and no limitations in the available quenches, i.e., $H_{n}=-\ln\rho_{X|n}$. Note that in the above expression we have incorporated the erasure cost, $k_{B}TI_{X;M}=k_{B}T[S(\rho_{0})-S(\rho_{X|A})/2-S(\rho_{X|B})/2]=k_{B}TS(p)=k_{B}T\log 2$. This is because, in contrast to the feedback control case, random quenching without observation requires no processing of information and hence a fair comparison requires taking that cost into account. The above equation tells us that despite having learned only one bit about $m$, the thermodynamic consequences for the distribution of the microstates $\mathbf{x}\in\mathbb{X}$ can be arbitrarily large.
@@ IV Optimal work extraction with control limitations
Importantly, rather than taking an arbitrary collection of Hamiltonians, in many situations we would be interested in constructing an optimal set of quenching Hamiltonians $\mathcal{H}$ with a limited dimension $N$, given an initial distribution $\rho_{0}(\mathbf{x})$ and given the conditional probability $p_{M|\mathbf{x}}(\mathbb{m})$ characterizing the noisy measurement of the system.
As mentioned before, the measurement results $\mathbf{m}$ can be generically distributed into $N$ disjoint but otherwise arbitrary sets $l(n)$ with $n=1,...,N$ to which one of the possible quenching Hamiltonians $H_{n}(\mathbf{x})\in\mathcal{H}$ is attached by using the lookup table defined by the minimization in Eq. (14). Constructing the optimal set $\mathcal{H}$ requires that the $H_{n}$ within each set $l(n)$ minimizes the expected entropy production across all such partition (see Fig. 1). The optimal form of $H_{n}(\mathbf{x})$ given a partition ${L}=\{\ell(n)\}_{n=1}^{N}$ is derived in Appendix A and reads:
@@
Figure 1: Illustration of the construction of an optimal quenching repertoire $\mathcal{H}$. We group all possible post-measurement distributions $\rho_{X|\mathbf{m}}$ in the simplex $\Delta(\mathbb{X}^{d_{S}})$ into $N$ sets according to a partition $\ell(n)$. The partition encodes a lookup table $n(\mathbf{m})$ from possible measurement results $\mathbf{m}$ (colored squares) in the space $\mathds{M}^{d_{M}}$ to the set of possible quenching actions. Measurement results $\mathbf{m}$ arise from the nonequilibrium system of interest (big grey square) by sampling from a noisy channel $p(\mathbf{m}|\mathbf{x})$, where system microstate $\mathbf{x}\in\mathbb{X}^{d_{S}}$ is distributed as $\rho_{0}$. The optimal quenches $\rho_{n}^{\rm opt}$ (large dots) within a given set $l(n)$ (green shaded area) is chosen by finding the distribution that minimizes the average entropy production (Kullback-Leibler divergence) when relaxing from all other post-measurement states within that set (small dots). Eq. 32 is used to obtain $H_{n}^{\mathrm{opt}}$ from $\rho_{n}^{\mathrm{opt}}$. Shaded regions depict a Voronoi-like partitioning of $\Delta(\mathbb{X}^{d_{S}})$ according to KL divergence from $\{\rho_{n}\}_{n=1}^{N}$. In the illustration, $N=4$, and the partition is $\{a,b\},\{c,d,e\},\{f\},\{g,h,i,j\}$.
Given the optimal form of the elements of the set $\mathcal{H}$ in Eq. (25), we can now compute the average entropy production (or work dissipated) in Eq. (12) for a given choice of partition ${L}$:
@@
$=\sum_{n}p(n)\left(S(\rho_{n}^{\rm opt})-\sum_{\mathbf{m}\in l(n)}p(\mathbf{m}|n)S(\rho_{X|\mathbf{m}})\right)$
@@
We note that in Eq. (27), the optimal quenching distribution $\rho_{n}^{\mathrm{opt}}(\mathbb{x})$ is a mixture of the post-measurement distributions $\rho_{X|\mathbf{m}}(\mathbb{x})$ across $\mathbf{m}$ weighted by $p(\mathbf{m}|n)$ [c.f. Eq. (25)], and that the parenthesized term in Eq. (27) is the mean entropy of $\rho_{X|\mathbf{m}}$ over the same mixing distribution $p(\mathbf{m}|n)$. As a consequence the entropy production can be rewritten as
@@
with $J_{\mathbf{p}}(\mathbf{P}):=S(Q)-\sum_{i}p_{i}S(P_{i})\geq 0$ denoting the generalized Jensen-Shannon divergence (JSD) among a weighted collection of distributions $\mathbf{P}=\{P_{i}\}_{i}$ with (normalized) weights $\mathbf{p}=\{p_{i}\}_{i}$ Lin (1991). Here $Q(\mathbb{x})=\sum_{i}p_{i}P_{i}(\mathbb{x})$ denotes the mixture distribution of $\{P_{i}(\mathbb{x})\}_{i}$ with weights $\{p_{i}\}_{i}$. (Note, $\sum_{i}p_{i}=1$ and $\sum_{\mathbb{x}}P_{i}(\mathbb{x})=1$.) In Eq. (28) we identified $P_{i}(\mathbb{x})=\rho_{X|\mathbb{m}}(\mathbb{x})$ and $p_{i}=p(\mathbb{m}|n)$ with $i=\mathbb{m}\in l(n)$, leading to $Q(\mathbb{x})=\rho_{n}^{\rm opt}(\mathbb{x})$, c.f. Eq (27). The Jensen-Shannon divergence and its generalizations have been employed in a variety of fields ranging from bioinformatics Sims et al. (2009) to the social sciences Klingenstein et al. (2014) and more recently in machine learning Englesson and Azizpour (2024). The above result in Eq. (28) implies that the optimal partition is one that minimizes the generalized JSD among the distributions grouped into a common partition block $\ell(n)$, averaged over the action $n$ taken.
A second, illuminating way of rewriting $\langle S_{\mathrm{tot}}\rangle$ in Eq. (27) is in terms of informational quantities. A key fact is that the distribution $\rho_{n}^{\rm opt}(\mathbb{x})$ in Eq. (26) is the marginal probability of finding the system in state $\mathbb{x}$ given that the measurement result is inside the partition $l(n)$. More precisely, we can rewrite the entropy production as the conditional mutual information Cover and Thomas (2012) between the system state and the measurement result, given the choice of quenching Hamiltonian according to the partitions $l(n)$:
$\langle S_{\mathrm{tot}}\rangle$ $=\sum_{n}p(n)\sum_{\mathbb{x}}\sum_{\mathbb{m}\in l(n)}p(\mathbb{m}|n)\rho_{X|\mathbb{m}}(\mathbb{x})\ln\left(\frac{\rho_{X|\mathbb{m}}(\mathbb{x})}{\rho_{n}^{\rm opt}(\mathbb{x})}\right)$
$=\sum_{n}p(n)\sum_{\mathbb{x}}\sum_{\mathbb{m}\in l(n)}P(\mathbb{x},\mathbb{m}|n)\ln\left(\frac{P(\mathbb{x},\mathbb{m}|n)}{\rho_{n}^{\rm opt}(\mathbb{x})p(\mathbb{m}|n)}\right)$
$=I_{X;M|{L}}\geq 0,$ (29)
where in the second line we expanded the joint conditional probability $P(\mathbb{x},\mathbb{m}|n)=p(\mathbb{m}|n)\rho_{X|\mathbb{m}}(\mathbb{x})$, and inside the logarithm we multiplied and divided by $p(\mathbb{m}|n)$.
Eq. (IV) says that achieving exactly zero entropy production requires that for all the partitions $l(n)$, there must be zero conditional mutual information $I_{X;M|{L}}$ between $\mathbb{x}$ and $\mathbb{m}$ given that the measurement result $\mathbf{m}$ is in the set $l(n)$. In other words, it requires the post-measurement distribution $\rho_{X|\mathbb{m}}(\mathbb{x})$ to actually be independent of the measurement result $\mathbb{m}$ within every partition $l(n)\in{L}$. Note that this effect arises only because of the restriction on the number of possible actions. Without such a restriction, this information-theoretic loss of the maximal amount of free energy the the agent can extract would not exist.
Notice though that such a condition cannot be satisfied without spoiling the possibility of performing feedback control, which in general requires correlations between $\mathbb{m}$ and $\mathbb{x}$, c.f. Eq. (11). In other words, zero entropy production requires that the measurements provide no other information than the one necessary to deduce the actual set $n$ to which the measurement belongs. In that case the distributions $\rho_{X|\mathbb{m}}(\mathbb{x})$ only depends on $n$, which would imply actually $\rho_{n}^{\rm opt}(\mathbb{x})=\rho_{X|\mathbb{m}}(\mathbb{x})$. That condition would be (trivially) satisfied when the number of possible quenching Hamiltonians coincide with the number of measurement results, $N=|\mathbb{M}^{d_{M}}|$, as expected.
By plugging Eq. (IV) into Eq. (11) we obtain the average extractable work:
$\langle{W}\rangle=$ $~{}k_{B}T\left[I_{X;M}-I_{X;M|{L}}\right]+\Delta\mathcal{F}_{H}.$ (30)
The term $I_{X;M}-I_{X;M|{L}}$ is sometimes called the “interaction information”. In general, it can be either positive or negative Cover and Thomas (2012). Therefore, an efficient measurement generating high correlations between $\mathbb{m}$ and $\mathbb{x}$ in general improves work extraction, in that it provides more information $I_{X;M}$ and in the limit of perfect correlation allows the extraction of all the free energy in the initial state. At the same time, spurious correlations as characterized by the conditional mutual information $I_{X;M|{L}}$ may spoil the ability to extract work by increasing the entropy production.
Taking into account the erasure of information after the feedback control protocol implies, nonetheless, that the optimal choice of partition corresponds to the one minimizing $I_{X;M|{L}}$. In other words, it corresponds to one in which the measurements generate the minimum mutual information between the measurement outcomes and the system state within the partitions $l(n)$, hence allowing the correct choice of the quenching Hamiltonians $H_{n}(\mathbb{x})=-k_{B}T\ln\rho_{n}^{\rm opt}(\mathbb{x})$ but avoiding, as much as possible, the correlations that can not be used in the feedback control operation.
Finally, we can straightforwardly obtain an upper bound on the entropy production as
$\sum_{n}p(n)S(\rho_{n}^{\rm opt})\geq\langle S_{\mathrm{tot}}\rangle,$ (31)
which follows from the concavity of entropy, $S(\rho_{n}^{\rm opt})\geq\sum_{\mathbb{m}\in l(n)}p(\mathbb{m}|n)S(\rho_{X|\mathbb{m}})\geq 0$. The above upper bound establishes respectively a lower bound on the average work that can be extracted, c.f. Eq. (11).
@@ V.1 Feedback control with memory limitations and noise-free measurements
As a standard scenario of general interest in stochastic thermodynamics we consider in the following the case in which the measurement results $m_{i}\in\mathbf{m}$ are associated one to one to the system variables $x_{i}\in\mathbf{x}$ and hence $\mathbb{M}=\mathbb{X}$ and $d_{X}=d_{M}$. Moreover, even if we are allowed to optimally choose the set of quenching Hamiltonians in the set $\mathcal{H}$, we assume operation with a finite memory where the number of quenches $N$ is generically much smaller than the state space of the system $N<|\mathbb{X}^{d_{X}}|$, and hence likewise smaller than the possible measurement results $|\mathbb{M}^{d_{M}}|=|\mathbb{X}^{d_{X}}|$.
@@
where $\mathds{1}\{\mathbf{x}\in\ell(n)\}$ is the indicator function within the partition $l(n)$, and $p^{\rm cg}_{0}(n):=\sum_{x\in\ell(n)}\rho_{0}(\mathbf{x})$ is the coarse-grained version of $\rho_{0}(\mathbf{x})$ according to partition $\ell(n)$. That is, the above quenching Hamiltonian takes finite values in the interval $\mathbf{x}\in l(n)$ and it is infinite otherwise. Here above the optimal post-quench distribution $\rho_{n}^{\rm opt}(\mathbf{x})$ is just the expression inside the parenthesis.
@@
$\langle S_{\mathrm{tot}}\rangle$ $=\sum_{\mathbf{m}}\rho_{0}(\mathbf{m})D(\delta_{\mathbf{m},\mathbf{x}}||e^{-\beta H_{n}})$
@@
Figure 2: Optimal partition $\ell(n)$ for a particle in one dimension using $N=8$ contiguous regions, $(x_{0},x_{1}]$, $(x_{1},x_{2}]$, …, $(x_{7},x_{8})$ according to Eq. (36). The solid line corresponds to the initial distribution, here taken to be $\rho_{0}(\mathbf{x})=e^{-x^{2}/2}/\sqrt{2\pi}$ while the blue dashed curve corresponds to the quenching Hamiltonian $H_{n}(\mathbf{x})$ in Eq. (32) for $n=2$ (not normalized, for visibility), equalling $\infty$ outside of $(x_{1},x_{2}]$.
[…]
In this case, the unique way to achieve optimal quenching is to set partition $\ell(n)$ as $N$ intervals in $\mathds{R}$ with boundaries $x_{k}$ for $k=0,1,2,...,N$ at
$x_{k}=F_{0}^{-1}\left(\frac{n}{N}\right),$ (36)
[…]
The optimal partition $\ell(n)$ is shown in Fig. 2 for $N=8$ possible quenches. The optimal quenches inherit the form of the initial distribution $\rho_{0}$ within each interval as illustrated with the blue dashed line in the second partition ($n=2$).